OpenStudy (vampirediaries):

15. What are the real & complex solutions of the polynomial equation?

OpenStudy (vampirediaries):

\[ x ^{4}-41x ^{2}= -400\]

OpenStudy (kainui):

Make the substitution x^2=y and solve it like you normally would, try it out. =)

OpenStudy (vampirediaries):

Do you mean to rewrite the equation like -41x^2= -x^4 -400 ? If not, could you elaborate please?

OpenStudy (kainui):

Yeah exactly like that, it's just nice to change the letter you're using to remind you that it's a different variable.

OpenStudy (kainui):

Like, x is just a letter, it doesn't have any holy significance or meaning at all. You could never use x and all math would remain exactly the same.